The Henderson family and the Brown family each used their sprinklers last summer. The Henderson family's sprinkler was used for 40 hours. The Brown family's sprinkler was used for 20 hours. There was a combined total output of 1500 L of water. What was the water output rate for each sprinkler if the sum of the two rates was 45 L per hour?
Answer by ikleyn(52831)

.
The Henderson family and the Brown family each used their sprinklers last summer.
The Henderson family's sprinkler was used for 40 hours.
The Brown family's sprinkler was used for 20 hours.
There was a combined total output of 1500 L of water.
What was the water output rate for each sprinkler if the sum of the two rates was 45 L per hour?
~~~~~~~~~~~~~~~~

 
Write equations as you read the problem


      H +   B =   45    liters per hour   (1)    (combined rate)

    40H + 20B = 1500    liters            (2)    (total output)


From equation (1)  B = 45-H.  Substitute it into equation (2)


    40H + 20*(45-H) = 1500

    40H + 900 - 20H = 1500

    40H - 20H = 1500 - 900

        20H   =     600

          H   =     600/20 = 30.


ANSWER. H output rate is 30 Liters per hour;  B output rate is 45-30 = 15 liters per hour.


CHECK.  40*30 + 20*15 = 1500 liters, total output.    ! Correct !

Solved.